OpenStudy (anonymous):

sqrt3-sqrt6/sqrt3+sqrt6

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

\[\sqrt{3} - \sqrt{6} / \sqrt{3} + \sqrt{6}\]